# PrisonBreak: Jailbreaking Large Language Models With at Most Twenty-Five Targeted Bit-flips

### v2

@@ 8.2 Behavioral Analysis
Mechanistic analysis. Recent work [34] identified a refusal feature in language models that determines if the model will deny a harmful query. This feature—the refusal direction—is computed as the average difference in activations computed for the last token of harmful (HarmBench) and benign queries (Alpaca). We compute the refusal direction for clean models following the procedure in Arditi et al. [34]. We then measure the directional agreement between the last token activations of each attack and the refusal direction on HarmBench, using cosine similarity. If the similarity is higher, the model is expressing the refusal feature, which should increase the probability of a refusal.
@@
The leftmost plot in Figure 4 compares the agreement across different attacks on Vicuna-13B-v1.5. Prompt-based attacks like GCG show less agreement, by reducing a model’s focus on harmful inputs and bypassing refusal by limiting its expression [34]. Ortho removes the direction entirely (indicated by the constant cosine similarity of zero), suppressing a model’s ability to refuse. Interestingly, our attack reduces the refusal expression less than the others, with cosine similarities closely aligning with the clean model. This indicates a strong presence of the refusal feature, which is surprising given that we show models jailbroken by PrisonBreak refuse at a lower rate. It does not control the refusal direction, thus functioning differently from Ortho.
Decomposing activations. To analyze activation differences in jailbroken models, we decompose them using a dimensionality reduction technique, UMAP [68]. We draw 100 random samples from Alpaca and HarmBench, and for each attack, obtain the activation at the last token position. We run UMAP on activations collected from the 200 samples, and reduce them into two components222We compare activations from the clean model and models compromised by Ortho and our attack. Empirically, these models are close to the clean one in loss space, indicating their activations do not differ drastically.
@@
The two remaining figures show visualizations for Qwen2-7B; we consider the last layer, where activations are most distinct and closest to the model’s output (see Appendix C for more layers). The clean model separates benign and harmful prompts into two distinct clusters. The harmful cluster may represent a region in the activation space where the model has learned to refuse. Prompt-based attacks like GCG produce activations that are similar for both harmful and benign inputs. This confirms that the adversarial suffix induces a distribution shift in the inputs, bypassing refusal by producing activations that differ from what the model can identify. Similarly, for Ortho, most activations from harmful inputs are near the benign ones, and align with the benign activations of the clean model. This suggests Ortho suppresses the refusal features, causing the jailbroken model to respond to all inputs. Our attack shows different behavior. The cluster for benign inputs overlaps significantly with that from the clean model, while harmful inputs form a distinct cluster. Our attack preserves a model’s utility but does not fully suppress the refusal features like Ortho, as shown in our mechanistic analysis. PrisonBreak may instead amplify activations associated with responding.
@@ 9 Potential Countermeasures
Potential countermeasures can defend against Rowhammer, jailbreaking, and bit-flips in the model. Many defenses have been proposed to mitigate Rowhammer [16, 69, 70], but increased latency, memory, and runtime overhead have prevented real-world deployment. Several works have also produced effective jailbreaking defenses [71, 72, 73, 74]; although designed primarily for prompt-based attacks, certain algorithms apply to other attack vectors, such as bit-flips. Model-based bit-flip defenses [75, 76, 77, 78] are related to our work, but many consider incompatible objectives or incur substantial computational overhead for commercial-scale language models. For these reasons, we focus on low-overhead and practically applicable jailbreaking- and model-based defenses, providing an in-depth overview of all defense paradigms in Appendix E.
@@
For model-based defenses, we examine Ranger [78], which clamps activations during inference, and quantized-weight reconstruction [79] which contains weights within bounds set by locally computed averages. Both defenses aim to mitigate the large parameter changes caused by bit-flip attacks. For Ranger, we compute bounds using 20% of Alpaca data (following the original study) and clip the outputs of each Transformer block. We adapt the defense by Li et al. [79] by clipping each layer’s weights to their original min. and max. values before each generation.
@@
For the jailbreaking-based defenses, we evaluate SafeDecoding [74], which fine-tunes a separate “safer” version of the victim model to steer logits away from harmful completions, and Intention Analysis [80], which prompts the model to reason about the user’s intent and potential harms before generating a response. For SafeDecoding, we utilize the publicly available implementation and fine-tuned models. For Intention Analysis, we apply the exact ‘one-pass’ prompt introduced in the original work to all queries.
@@
In Table VIII, we show that our attack remains effective against these defenses. Ranger is ineffective, yielding no reduction in ASR. This is likely because the large activations typical in language models [81] produce clamping bounds that are not exceeded, even after substantial weight value changes. Weight clipping (WC) effectively reduces ASR from 87.2% to 9.6%, comparable to clean models (9.4%). But, against an adaptive adversary (Adaptive) who selectively inflicts bit-flips while keeping parameter changes within specified layer-wise bounds, the attack regains effectiveness (70.2%). SafeDecoding (SD) is less effective, reducing ASR to 41.5%, though its impact varies with the robustness of the underlying safety-tuned models. An adaptive attacker targeting these fine-tuned models can restore ASR to 86.0%, nearly matching the undefended baseline. Intention Analysis (IA) is the most effective, reducing ASR to 7.8%, lower than that observed in clean models. The reformatted prompt significantly shifts the output distribution, rendering original bit-flips ineffective. However, an adaptive adversary can identify bit-flips that remain effective under IA’s reformatted prompts, recovering ASR to 88.7%.
@@ 10 Conclusion
Our work exposes a new vulnerability in commercial-scale language models: the ability of these models to refuse answering harmful queries, a feature learned through safety alignment mechanisms, can be broken by manipulating only a few parameters out of billions. This finding is surprising and in stark contrast to prior work [34, 33], which requires altering millions of parameters to remove the refusal ability. We introduce PrisonBreak, a new attack that exploits this vulnerability. The adversary induces targeted bit-flips at memory locations storing critical parameters to jailbreak language models at runtime. We also present an efficient offline procedure for identifying target bits, which reduces the computations required to locate critical bits by over an order of magnitude compared to prior attacks. The attacker then leverages Rowhammer, a software-induced fault attack, to flip those target bits in memory and jailbreak the model. In our evaluation of 10 open-source language models, we show that the attacker can PrisonBreak these models with as few as 5 and no more than 25 bit-flips. Our jailbroken models generate high-quality responses to harmful queries with an attack success of 80–98% while preserving their utility. We also demonstrate that an attacker can jailbreak target models without access to their parameters—identified bit-flips transfer between models that share architecture and pre-training. Our attack reliably jailbreaks all evaluated models even on systems targeted by prior bit-flip attacks. Moreover, our hardware-aware search algorithm can achieve successful jailbreaking on models deployed on highly Rowhammer-secure systems. We characterize the mechanisms behind our attack, and show PrisonBreak operates differently than existing jailbreaking attacks. We also discuss potential countermeasures to mitigate the risk.
@@ Acknowledgments
Zachary Coalson and Sanghyun Hong are supported by the Google Faculty Research Award. Jeonghyun Woo and Prashant J. Nair are supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) (#RGPIN-2019-05059) and a Gift from Meta Inc. Yu Sun and Lishan Yang are partially supported by the National Science Foundation (NSF) grants (#2402940 and #2410856) and the Commonwealth Cyber Initiative (CCI) grant (#HC‐3Q24‐047). Bo Fang is supported by the U.S. DOE Office of Science, Office of Advanced Scientific Computing Research, under award 66150: “CENATE - Center for Advanced Architecture Evaluation” project. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under contract DE-AC05-76RL01830. Any opinions, findings, and conclusions or recommendations expressed here are those of the authors and do not necessarily reflect the views of the funding agencies.
@@ Appendix A Experimental Setup in Detail
We implement our attack using Python v3.10.13 and PyTorch v2.4.1, with CUDA 12.1 for GPU usage. All language models and datasets used in our work are open-source and available on HuggingFace [55] or their respective repositories. Information regarding the DRAM modules used in our study is available from the original work [39]. We run the PrisonBreak evaluation on a machine with an Intel Xeon Processor with 48 cores, 64 GB of memory, and 8 Nvidia A40 GPUs. For our fault injection simulations, we use an Intel Xeon Processor with 40 cores and 128GB of memory.
Appendix B Impact of Attack Configurations
Here, we vary the attack configuration choices for an additional model: Vicuna-7b-v1.3. Results are presented in Figure 5. For the proxy dataset size, using just 5 samples is sufficient to achieve a high ASR (87.4%). We also find that increasing the number of candidates from 100150 lowers ASR from 94.3491.8%. These results differ from §6.5, where we found that a dataset size of is ineffective and setting substantially improved ASR on Llama2-7b. It is likely that because lesser-aligned models like Vicuna-7b-v1.3 are easier to jailbreak, PrisonBreak requires less data and fewer candidates. We observe largely similar results for the utility score threshold, but find the most aggressive value () less effective, as increasing it to raises ASR from 85.594.3% with negligible impact on Acc. We find that the tighter threshold required the analysis of substantially more candidates, meaning the selected bit-flips minimized the jailbreaking objective less effectively. Like §6.5, a manual analysis of the harmful generations finds that higher thresholds () result in lower fidelity outputs regardless of the reported Acc. on tinyBenchmarks. Finally, the rightmost figure shows that only 5 bit-flips are needed to achieve the best ASR, after which it saturates, supporting our finding in §6.2 that less-aligned models require fewer bit-flips to jailbreak.
Appendix C Activation Decomposition
@@
Appendix D Results on tinyBenchmarks
Table IX shows our full evaluation results for PrisonBreak on tinyBenchmarks. Notably, the change in performance differs across tasks. Models on average perform the worst on GSM8k—a benchmark consisting of grade school math problems—with an accuracy reduction of 5.61%. This may arise from the lack of these problems in the dataset we use to compute the utility score. The second most affected task is TruthfulQA with an average accuracy reduction of 2.75%. This fact is less surprising as jailbroken models are more likely to be deceptive [33, 34]. Other tasks exhibit smaller drops in accuracy; in general, we find that PrisonBreak preserves utility on benign tasks.
Module RV (ppm) Vicuna- 7b-v1.3 Vicuna- 7b-v1.5 Vicuna- 13b-v1.5 Llama2-7b Llama2-13b Llama3-8b Llama3-70b Tulu3-8b Qwen2-1.5b Qwen2-7b A17 64.5 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A2 62.7 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A5 52.7 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A16 46.4 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A6 45.8 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A8 43.1 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A0 38.3 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A18 37.5 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A14 29.8 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A13 27.8 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A15 24.5 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A12 20.3 100% 100% 100% 100% 100% 100% 100% 100% 100% 100% A7 14.9 99.9% 99.1% 99.7% 54.4% 99.6% 99.1% 99.1% 99.3% 99.1% 98.7% A1 5.7 94.5% 47.6% 63.3% 0.2% 65.9% 54.3% 57.3% 54.8% 46.1% 44.2% A19 5.4 99.6% 93.5% 97% 19.4% 96.9% 94.7% 95.1% 95.4% 93.5% 93.8% D0 5.0 91.7% 35.9% 52.4% 0% 55.5% 44.8% 46.8% 39.9% 32.6% 34.1% D3 3.2 80.2% 8.7% 16.7% 0% 19.5% 10.2% 15.2% 11.2% 4.9% 6.2% D1 3.1 96.30% 59.8% 76% 0.4% 74.8% 63.2% 68.2% 66.3% 56.1% 53.5% A4 2.4 22.50% 0% 0% 0% 0% 0% 0% 0% 0% 0% A9 2.3 64.00% 1.5% 4.4% 0% 4.4% 1.8% 3.1% 2.5% 0.6% 0.7% A11 1.5 38.10% 0% 0.4% 0% 0.4% 0.1% 0.1% 0% 0% 0% A10 1.4 35.20% 0% 0.2% 0% 0.1% 0% 0.4% 0% 0.1% 0% C2 1.2 23.90% 0% 0% 0% 0% 0% 0% 0% 0% 0% D2 0.9453 51.20% 0.2% 0.4% 0% 1% 0.3% 0.6% 0.5% 0% 0.1% A3 0.8130 10.70% 0% 0% 0% 0% 0% 0% 0% 0% 0% B9 0.5695 3.70% 0% 0% 0% 0% 0% 0% 0% 0% 0% B4 0.5155 1.90% 0% 0% 0% 0% 0% 0% 0% 0% 0% C5 0.4787 0.20% 0% 0% 0% 0% 0% 0% 0% 0% 0% C4 0.3581 0.70% 0% 0% 0% 0% 0% 0% 0% 0% 0% C3 0.2962 0.10% 0% 0% 0% 0% 0% 0% 0% 0% 0% B1 0.2356 0.10% 0% 0% 0% 0% 0% 0% 0% 0% 0% B8 0.1201 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B3 0.0517 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B6 0.0363 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B7 0.0326 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B0 0.0293 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% C1 0.0130 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% C0 0.0121 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B2 0.0070 0% 0% 0% 0% 0% 0% 0% 0% 0% 0% B5 0.0065 0% 0% 0% 0% 0% 0% 0% 0% 0% 0%
Appendix E Detailed Discussion on Countermeasures

### v3

@@ 8.2 Behavioral Analysis
Mechanistic analysis. Recent work [4] identified a refusal feature in language models that determines if the model will deny a query. This feature—the refusal direction—is computed as the average difference in activations computed for the last token of harmful (HarmBench) and benign (Alpaca) queries. We compute the refusal direction for clean models following the procedure in Arditi et al. [4]. We then measure the directional agreement between the last token activations of each attack and the refusal direction on HarmBench, using cosine similarity. If the similarity is higher, the model is expressing the refusal feature, which should increase refusal.
@@
The leftmost plot in Figure 4 compares the agreement across different attacks on Vicuna-13B-v1.5. Prompt-based attacks like GCG show less agreement; by reducing a model’s focus on harmful inputs, they bypass refusal by limiting its expression [4]. Ortho removes the direction entirely (indicated by the constant cosine similarity of zero), suppressing a model’s ability to refuse. Interestingly, our attack reduces the refusal expression less than the others, with cosine similarities closely aligning with the clean model. This indicates a strong presence of the refusal feature, which is surprising given that we show models jailbroken by PrisonBreak refuse at a lower rate. It does not control the refusal direction, thus functioning differently from Ortho.
@@
Decomposing activations. To analyze activation differences in jailbroken models, we decompose them using UMAP [56]. We draw 100 random samples from Alpaca and HarmBench, and extract the last-token activations for each of them. We run UMAP on these 200 activations,222We compare activations from the clean model (no attack) and models compromised by GCG, Ortho and our attack. Empirically, the compromised models are largely close to the clean model (no attack) in the loss space, indicating their activations do not differ drastically. and reduce them into two components. Figure 4 (center and right) shows these activation visualizations for Qwen2-7B; we consider the last layer, where activations are most distinct and closest to the model’s output (see Appendix D for more layers). The clean model (no attack) separates benign and harmful prompts into two distinct clusters; the harmful cluster may represent a region where the model has learned to refuse. Prompt-based attacks like GCG produce activations that are similar for both harmful and benign inputs. This confirms that the adversarial suffix induces a distribution shift in the inputs, bypassing refusal by producing activations that differ from what the model can identify. Similarly, for Ortho, most activations from harmful inputs are near the benign ones, and align with the benign activations of the clean model. This suggests Ortho suppresses the refusal features, causing the jailbroken model to respond to all inputs. Our attack shows different behavior. The cluster for benign inputs overlaps significantly with that from the clean model, while harmful inputs form a distinct cluster. Our attack preserves a model’s utility but does not fully suppress the refusal features like Ortho, as shown in our mechanistic analysis. It is likely that PrisonBreak may instead be amplifying activations associated with responding.
@@ 9 Potential Countermeasures
Potential countermeasures can defend against Rowhammer, jailbreaking, and bit-flips in the model. Many defenses have been proposed to mitigate Rowhammer [42, 41, 83], but increased latency, memory, and runtime overhead have prevented widespread deployment. Several works have also produced effective jailbreaking defenses [100, 96, 40, 90]; although designed primarily for prompt-based attacks, certain algorithms apply to other attack vectors, such as bit-flips. Model-based bit-flip defenses [52, 48, 85, 16] are related to our work, but many consider incompatible objectives or incur substantial computational overhead for commercial-scale LLMs. For these reasons, we focus on low-overhead and practically applicable jailbreaking- and model-based defenses, providing an in-depth overview of all paradigms in Appendix G.
@@
For model-based defenses, we examine Ranger [16], which clamps activations during inference, and quantized-weight reconstruction [49] which contains weights within bounds set by locally computed averages. Both defenses aim to mitigate large bit-flip-induced parameter changes. For Ranger, we compute bounds using 20% of Alpaca data and clip the outputs of each Transformer block. We adapt the defense by Li et al. [49] by clipping each layer’s weights to their original min. and max. values before each generation.
@@
For jailbreaking-based defenses, we evaluate SafeDecoding [90], which fine-tunes a “safer” version of the victim model to promote safer logits, and Intention Analysis [97], which prompts models to reason about the user’s intent before generating a response. For SafeDecoding, we utilize the open-source implementation and models; for Intention Analysis, we apply their exact ‘one-pass’ prompt to all queries.
@@
Table 7 shows that our attack remains effective against these defenses. Ranger is ineffective, yielding no reduction in ASR. This is likely because the large activations typical in language models [77] produce clamping bounds that are not exceeded, even after large weight changes. Weight clipping (WC) effectively reduces ASR from 87.2% to 9.6%, comparable to clean models (9.4%). But, with an adaptive adversary (Adaptive) who selectively inflicts bit-flips while keeping parameter changes within specified layer-wise bounds, the attack regains effectiveness (70.2%). SafeDecoding (SD) is less effective, reducing ASR to 41.5%, though its impact varies with the robustness of the underlying safety-tuned models. An adaptive attacker targeting these fine-tuned models can restore ASR to 86.0%, nearly matching the undefended baseline. Intention Analysis (IA) is the most effective, reducing ASR to 7.8%, lower than that observed in clean models. The reformatted prompt significantly shifts the output distribution, rendering original bit-flips ineffective. However, an adaptive adversary can identify bit-flips that remain effective under IA’s reformatted prompts, recovering ASR to 88.7%.
@@ 10 Conclusion
We expose a critical vulnerability in safety-aligned LLMs: their refusal to answer harmful queries can be broken by bit-flipping a handful of parameters. Our attack, PrisonBreak, efficiently identifies and flips critical bits in the model, enabling jailbreaks via Rowhammer with 5–25 bit-flips, orders of magnitude fewer than prior methods [4, 93]. Across 5 open-source model families (Vicuna, Llama2, Llama3, Tulu3, Qwen2), PrisonBreak achieves a jailbreaking ASR of 69–91% with Rowhammer on a GDDR6 GPU and 80-98% in a hardware-agnostic attack, while preserving model utility. We also discuss countermeasures to mitigate the risk.
@@ Acknowledgments
Zachary Coalson and Sanghyun Hong are supported by the Google Faculty Research Award 2023. Jeonghyun Woo and Prashant J. Nair are supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) (#RGPIN-2019-05059) and a Gift from Meta Inc. Chris S. Lin, Joyce Qu and Gururaj Sailshwar are supported by NSERC (#RGPIN-2023-04796), and an NSERC-CSE Research Communities Grant under funding reference number ALLRP-588144-23. Yu Sun and Lishan Yang are partially supported by the National Science Foundation (NSF) grants (#2402940 and #2410856) and the Commonwealth Cyber Initiative (CCI) grant (#HC‐3Q24‐047). Bo Fang is supported by the U.S. DOE Office of Science, Office of Advanced Scientific Computing Research, under award 66150: “CENATE - Center for Advanced Architecture Evaluation” project. The Pacific Northwest National Laboratory is operated by Battelle for the U.S. Department of Energy under contract DE-AC05-76RL01830. Any opinions, findings, and conclusions or recommendations expressed here are those of the authors and do not necessarily reflect the views of the funding agencies.
@@ Appendix A Experimental Setup in Detail
We implement our attack using Python v3.10.13 and PyTorch v2.4.1, with CUDA 12.1 for GPU usage. All language models and datasets used in our work are open-source and available on HuggingFace [88] or their respective repositories. We run the PrisonBreak (hardware-agnostic) evaluation on a machine with an Intel Xeon Processor with 48 cores, 64 GB of memory, and 8 Nvidia A40 GPUs. Our end-to-end Rowhammer experiments are performed on a machine with an AMD Ryzen processor with 12 cores and 32 GB of memory, and an RTX A6000 GPU with 48 GB of GDDR6 DRAM.
Appendix B Hammering Campaign Results
DRAM Bank Flip Direction Bit Location in Byte Location in float16 Weight A 4 M4 B 0 M0 B 6 M6 B 7 M7 C 0 M8 D 6 E4 D 0 M8 D 6 E4 G 6 M6 H 4 E2 E = Exponent, M = Mantissa, (E/M)x = xth bit in E/M
Table 8 reports the locations of the ten bit-flips identified in our offline hammering campaign on the NVIDIA A6000, spanning eight banks (A–H). Most bit-flips occur in even bytes, which map to Mantissa bits (Mx) of half-precision floating-point weights and therefore induce only small parameter changes. Notably, only two bit-flips map to the most significant exponent bit (E4), which we use exclusively for high jailbreaking success in the end-to-end exploitation (§7).
Appendix C tinyBenchmarks Results
Table 9 shows our full evaluation results for PrisonBreak on tinyBenchmarks. Notably, the change in performance differs across tasks. Models on average perform the worst on GSM8k—a benchmark consisting of grade school math problems—with an accuracy reduction of 5.6%. This may arise from the lack of these problems in the dataset we use to compute the utility score. The second most affected task is TruthfulQA with an average accuracy reduction of 2.8%. This fact is less surprising as jailbroken models are more likely to be deceptive [93, 4]. Other tasks exhibit smaller drops in accuracy; in general, we find that PrisonBreak preserves utility on benign tasks.
Appendix D Activation Decomposition
@@
Appendix E Impact of Model Datatype
We study how the model’s datatype (the numerical format of its weights) influences jailbreaking success. We evaluate four widely used datatypes supported by PyTorch: full-precision floating point (float32), half-precision floating point (float16), brain floating point (bfloat16), and 8-bit integer quantization (int8) [21]. Since PyTorch’s int8 implementation does not support gradient computations, we compute gradients for weight ranking by de-quantizing the model at each epoch, while still applying bit flips to the original int8 version. For comparability, we search over all exponent bits and the sign bit for floating-point formats; for int8, we find that only the three most significant bits induce substantial changes and therefore only target them. We use for all experiments. Figure 6 shows the ASR and Acc. for each datatype on Vicuna-7B-v1.3.
PrisonBreak is effective against all floating-point datatypes, achieving ASRs from 89.3–93.1%. The float32 datatype appears slightly more vulnerable, likely because its larger exponent size enables more precise weight modifications. In contrast, the quantized int8 model is more resilient, with an ASR of only 36.5%. This aligns with prior work showing that quantized models are generally more robust to BFAs [33, 94]. Nevertheless, our attack still improves ASR by 24.5% points compared to the clean model, and additional flips may further increase success. We leave further exploration of jailbreaking quantized models to future work. Across datatypes, the attack has minimal effect on model utility: Acc. remains between 48.3–49.7%, within 2–3% of the base model. Overall, PrisonBreak is versatile with respect to the weight representation of victim LLMs.
Appendix F Impact of Attack Configurations
Here, we vary the attack configuration choices for an additional model: Vicuna-7b-v1.3. Results are presented in Figure 7. For the proxy dataset size, using just 5 samples is sufficient to achieve a high ASR (87.4%). We also find that increasing the number of candidates from 100150 lowers ASR from 94.391.8%. These results differ from §6.5, where we found that a dataset size of is ineffective and setting substantially improved ASR on Llama2-7b. It is likely that because lesser-aligned models like Vicuna-7b-v1.3 are easier to jailbreak, PrisonBreak requires less data and fewer candidates. We observe largely similar results for the utility score threshold, but find the most aggressive value () less effective, as increasing it to raises ASR from 85.594.3% with negligible impact on Acc. We find that the tighter threshold required the analysis of substantially more candidates, meaning the selected bit-flips minimized the jailbreaking objective less effectively. Like §6.5, a manual analysis of the harmful generations finds that higher thresholds () result in lower fidelity outputs regardless of the reported Acc. on tinyBenchmarks. Finally, the rightmost figure shows that only 5 bit-flips are needed to achieve the best ASR, after which it saturates, supporting our finding in §6.2 that less-aligned models require fewer bit-flips to jailbreak.
Appendix G Detailed Discussion on Countermeasures